The perimeter of a rectangle is 304 cm the length is 40 cm longer than the width what is the length and width

Answer by ankor@dixie-net.com(22740)

The perimeter of a rectangle is 304 cm
2L + 2w = 304
simplify, divide by 2
L + w = 152
:
the length is 40 cm longer than the width
L = w + 40
:
what is the length and width
In the first simplified equation, replace L with (w+40)
(w+40) + w = 152
2w = 152 - 40
2w = 112
w = 112/2
w = 56 cm is the width
then
L = 56 + 40
L = 96 cm is length
:
:
Check this, find the perimeter using these solutions
2(96) + 2(56) =
192 + 112 = 304